It takes Ricky, traveling at
36 mph,
30 minutes longer to go a certain distance than it takes Ross traveling at
63 mph. Find the distance traveled.

Answer by stanbon(75887)

It takes Ricky, traveling at 36 mph, 30 minutes longer to go a certain distance than it takes Ross traveling at 63 mph. Find the distance traveled.
-----
Ricky DATA:
rate = 36 mph ; time = t + 1/2 hrs ; distance = 36t + 18 miles
------
Ross DATA:
rate = 63 mph ; time = t hrs ; distance = 63t miles
-------
Equation::
distance = distance
36t + 18 = 63t
27t = 18
t = 18/27 = 2/3 hr
-----
Ans: distance = 63t = 63(2/3) = 42 miles
---------------
Cheers,
Stan H.
--------------